# Local asymptotics of selection models with applications in Bayesian selective inference

### v2

@@ 3 Asymptotic expansion of the likelihood
First, we examine how selection acts on Gaussian models. Let $Z\sim N(h,\Sigma)$, with parameter $h\in\mathbb{R}^{p}$ and fixed positive-definite covariance $\Sigma\in\mathbb{R}^{p\times p}$, and consider the selection model induced by an arbitrary selection function $p(z)$. The selective log-likelihood ratio for this model around $h=0$ is given by
$\log\frac{f_{h}(Z)}{f_{0}(Z)}=h^{T}\Sigma^{-1}Z-\frac{1}{2}h^{T}\Sigma^{-1}h+\log\frac{\varphi(0)}{\varphi(h)},\quad\varphi(h)=E_{h}[p(Z)].$ (3)
@@
The main result of this paper, Theorem 1, establishes conditions under which the sequence of selective log-likelihood ratios of a given selection model $\mathcal{M}_{n}(u_{n})$ admits an asymptotic expansion of the form (3) for a certain specification of selection function, which in general depends on the true generating parameter $\theta$. It is worth remarking that, since selection functions can vary arbitrarily with $n$, $\{\mathcal{M}_{n}(u_{n})\colon n\in\mathbb{N}\}$ does not in general converge to a fixed Gaussian selection model. Instead, it can, under suitable conditions, be approximated by a sequence of such. Before stating the Theorem, we need to introduce some notation.
Notation. In the non-selective model $\mathcal{M}_{n}$, let $l_{\theta}(y_{i})=\log f_{\theta}(y_{i})$ denote the log-likelihood for the $i$-th observation, $\triangledown l_{\theta}(y_{i})$ the score, and $I_{\theta}=E_{\theta}[\triangledown l_{\theta}(Y_{i})\triangledown l_{\theta}(Y_{i})^{T}]$ the per-observation Fisher information, assumed to be positive-definite for all $\theta\in\Theta$. For a function $g:I\subseteq\mathbb{R}^{p}\to\mathbb{R}$, consider its Lipschitz norm,
@@
[…]
@@
Suppose that $\{F_{\theta}\colon\theta\in\Theta\}$ is DQM, let $Z_{n}\equiv Z_{n}(\theta)=n^{-1/2}I_{\theta}^{-1}\sum_{i=1}^{n}\triangledown l_{\theta}(Y_{i})$, and assume that the sequence of selection functions is uniformly bounded:
@@
For a fixed $\theta\in\Theta$, assume that $M(x)=E_{\theta}[\exp\{x^{T}\triangledown l_{\theta}(Y_{1})\}]<\infty$ in a neighborhood of the origin, that there exists a sequence of measurable functions $\{p^{*}_{n}(\cdot;\theta)\colon\mathbb{R}^{p}\to[0,\infty)\colon n\in\mathbb{N}\}$ satisfying $p^{*}_{n}(z;\theta)=E_{\theta}\left[p_{n}(Y^{n})\mid Z_{n}=z\right]$ for all $z$ in the support of $Z_{n}$, and that either of the following conditions are met:
@@
$\log\frac{f_{\theta+h_{n}/\sqrt{n}}(Y^{n}\mid u_{n})}{f_{\theta}(Y^{n}\mid u_{n})}=h^{T}I_{\theta}Z_{n}-\frac{1}{2}h^{T}I_{\theta}h+\log\frac{\varphi_{n}^{*}(0)}{\varphi_{n}^{*}(h)}+o_{F^{u_{n}}_{\theta}}(1).$ (8)
@@
The proof of Theorem 1 can be found in the supplementary material along with the proofs of all the other theoretical results of the paper. It takes as starting point the LAN property of the underlying, non-selective model, from which the first two terms of the expansion can be derived. The main technical difficulty lies in validating the approximation of the probability ratios $\varphi_{n}(\theta+n^{-1/2}h)/\varphi_{n}(\theta)$ by their Gaussian counterparts. This requires consideration of the moment generating function of $Z_{n}$ given selection. The existence of $M(x)$, together with the regularity assumptions on the selection functions in the limiting Gaussian models, prove to be sufficient to ensure uniform validity. In addition, the conditions on the selection functions are fairly minimal, and satisfied in most standard applications of the framework, as demonstrated in the examples considered below.
The requirement that $\varphi_{n}(\theta)$ is bounded away from zero is needed to ensure that the Gaussian approximation of the normalized score remains valid after conditioning on selection and to preserve the scale of likelihood ratios. This condition guarantees asymptotic regularity of the model even if selection is deterministic, that is, when the selection functions are of the form $p_{n}(y^{n})=\mathbf{1}(y^{n}\in A_{n})$ for some events $A_{n}$. However, it is well-known that, in these scenarios, if $y^{n}$ falls close to the boundary of $A_{n}$, irregularities can arise, leading to poor inferential performance unless $n$ is very large.
Scenarios with $\varphi_{n}(\theta)\to 0$ can arise in models involving many parameters. While this asymptotic regime is not covered by the Theorem, in some scenarios it is possible to verify the boundedness condition either exactly or asymptotically:
@@
    If the selection condition is rejection of a simple hypothesis $H_{0}\colon\theta=\theta_{0}$ at a level $\alpha$, then $\varphi_{n}(\theta)\geq\alpha$ for all $n$, provided the test is consistent. Thus, as long as $\alpha$ does vanish as $n\to\infty$, the condition is satisfied. This is typically the case for multiple comparisons models, such as the one described in Section 3.2, with a bounded number of parameters.
@@
Selection rules that do not allow for the existence of $\sqrt{n}$-consistent estimators are typically deterministic and yield inferences with very limited power. Their lack of power often requires the collection of additional data to draw meaningful conclusions about the parameter, leading to a data carving scenario in which estimation of $\theta$ with the required accuracy is feasible.
It is also important to note that the selection function in the asymptotic model, $p^{*}_{n}(z;\theta)$, depends on the true parameter $\theta$. This dependency introduces additional complexity to the asymptotic analysis. Under the LAN framework, the true value of $\theta$ only influences the covariance of the limiting Gaussian location model, leaving the group structure unaltered and facilitating a unified approach to analyzing the properties of inferential procedures. By contrast, under selection, $\theta$ has a deeper structural influence on the limiting model through its presence in $p^{*}_{n}$. This may affect the asymptotic distribution of estimators and test functions, as well as optimal choice of inferential approach. It remains to be explored whether the true asymptotic selection function can be effectively approximated by $\hat{p}^{*}_{n}(z)=p^{*}_{n}(z;\hat{\theta})$, where $\hat{\theta}$ is a $\sqrt{n}$-consistent estimator of $\theta$.
The following examples are representative of many common situations encountered in selective inference.
@@
[Deterministic selection] Let $Y_{1},\ldots,Y_{n}$ be an IID sample from an exponential family distribution with density
$f_{\theta}(y_{i})=h(y_{i})\exp\left\{\theta^{T}y_{i}-A(\theta)\right\},\quad\theta\in\Theta.$ (9)
@@
[Data carving] Consider the same setting as in the previous example, but with selection condition given by $\bar{Y}^{(1)}_{n}\in E_{n}$, where $\bar{Y}^{(1)}_{n}$ is the mean of the first $n/2$ samples, where $n$ is assumed to be even. Then,
@@
Condition 2 of the Theorem is satisfied provided $\sup\{\|\log f_{\bar{Y}^{(1)}_{n}}\|_{\text{L}}\colon n\geq 1\}<\infty$. Conditioning on selection based on a subsample of the data was introduced by Fithian et al. (2017) and is commonly referred to as data carving.
@@
[Randomization] Consider the same setting, but suppose now that the selection mechanism acts on $U_{n}=\sqrt{n}\bar{Y}_{n}+W$, where the noise $W$ has a known density $f_{W}(w)$, so that selection takes the form $U_{n}\in E_{n}$. Then, $p^{*}_{n}(z;\theta)=P\{\sqrt{n}\bar{y}_{n}(z)+W\in E_{n}\}$, so
@@
which is bounded provided $f_{W}(w)$ is bounded. Proposed by Tian and Taylor (2018), randomized selection has become common in practice in post-selection inference, as it enables more powerful analyses.
@@
[Condition on randomized statistic] Consider the same setting as in the previous example. In cases where knowledge of the precise form of $E_{n}$ is unavailable to the statistician, it is common to condition on the observed value of $U_{n}$ instead (García Rasines and Young, 2023), and the selection function becomes $p^{*}_{n}(z;\theta)=f_{W}\{u-\triangledown^{2}A(\theta)z+\sqrt{n}\triangledown A(\theta)\}$, which satisfies the required regularity conditions for most standard noise distributions, such as the standard Gaussian or Laplace distributions (Tian and Taylor, 2018).
3.1 A simple univariate example
Let us assume that the underlying model is exponential of rate parameter $\theta$, and that for scientific reasons we are interested in large values of $\theta$. That is, we would like to perform inference on $\theta$ only if the sample average, $\bar{Y}_{n}$, is small, indicating a large rate. We will illustrate the four previous types of selection mechanism in this context.
A deterministic selection mechanism would be of the form $\bar{Y}_{n}<t_{n}$ for some predetermined threshold value $t_{n}$, which might have been obtained through a power analysis or in some other data-independent way. For a given $\theta$ and realization $\bar{Y}_{n}=\bar{y}_{n}$, the normalized score for this model is $z=\theta^{2}\sqrt{n}(1/\theta-\bar{y}_{n})$, and the selection function in the asymptotic model becomes $p^{*}_{n}(z;\theta)=\mathbf{1}\{z>\theta^{2}\sqrt{n}(1/\theta-t_{n})\}$. If the analogous selection rule is applied only to half the samples, so that inference is performed whenever $\bar{Y}_{n}^{(1)}<t_{n}$, then, assuming $n$ is even and noticing that $Z_{n}$ follows a shifted and scaled gamma distribution, we obtain, after some manipulation,
@@ 3.1 A simple univariate example
Finally, suppose selection acts on $U_{n}=\sqrt{n}\bar{Y}_{n}+W$, where $W\sim N(0,\sigma^{2}_{W})$ for some pre-determined $\sigma_{W}>0$, so that the sample gets selected if $U<t_{n}$. Then,
$p^{*}_{n}(z;\theta)=\Phi\left\{\frac{\sqrt{n}}{\sigma_{W}}\left(t_{n}-\frac{1}{\theta}\right)+\frac{z}{\sigma_{W}\theta^{2}}\right\},$
where $\Phi$ is the $N(0,1)$ distribution function. If, instead, we condition on the observed value of $U_{n}$, $u_{n}$ (perhaps due to unavailability of knowledge of the precise threshold $t_{n}$ applied), the selection function becomes
$p^{*}_{n}(z;\theta)=\phi\left\{\frac{1}{\sigma_{W}}\left(u_{n}-\frac{\sqrt{n}}{\theta}\right)+\frac{z}{\theta^{2}}\right\},$
with $\phi$ the $N(0,1)$ density function.
Figure 1 shows $r_{n}(h;\theta)=\log\{\varphi_{n}(\theta+h/\sqrt{n})/\varphi_{n}(\theta)\}$ and $r^{*}_{n}(h;\theta)=\log\{\varphi^{*}_{n}(h)/\varphi^{*}_{n}(0)\}$ in the four situations considered above, for $\theta=2$, $n=40$, and $h\in[-2,2]$. The main step in the proof of Theorem 1 is to establish that the two functions approximate each other uniformly in compact sets of $h$ as $n\to\infty$. This figure provides visual evidence of the approximation across the range of values of $h$ considered. In all cases, the functions $\varphi^{*}_{n}(h)$ where computed by numerical integration of $I_{\theta}^{1/2}\phi(I_{\theta}^{1/2}z)p^{*}_{n}(z;\theta)$ over $z$.
@@ 3.2 Inference on winners
In this section we explore a classic example of selection bias, where inference is conducted on the most promising parameters chosen from a large set of candidates. This constitutes a paradigmatic application of selective inference in the IID setting. In such problems, $N$ independent populations under investigation, with respective parameters $\theta_{1},\ldots,\theta_{N}$, are ranked based on a data-dependent criterion (typically some measure of significance of the parameter), and only the top-ranked populations are selected for further analysis. For example, we may compute $N$ $p$-values of significance and select for formal inference only those parameters whose respective $p$-values are among the smallest $K<N$ or fall below a fixed significance threshold $t$. Once a subset of parameters has been selected, further data may be collected on the selected populations, leading to a data-carving scenario, as described in the previous section. This type of situation is common in clinical studies, where multiple treatments are tested and only the most promising ones are advanced for further investigation. Another common application occurs in genomics, where a vast number of gene expressions are measured, and only those showing the strongest signals, such as significant associations with a disease, are formally analyzed.
@@
We will consider the following specification of the problem. For $j\in\{1,\ldots,N\}$ and $i\in\{1,\ldots,n_{1}\}$, let $Y_{i}^{j}\sim N(\mu_{j},\sigma^{2}_{j})$ independently, where all $\mu_{j}$’s and $\sigma^{2}_{j}$’s are unknown. That is, we have random samples of size $n_{1}$ from $N$ different Gaussian populations. Suppose that we are only interested in means with large values. Accordingly, we compute $N$ $t$-statistics, $\{S_{j}=\bar{Y}^{j}/\sqrt{V^{j}/n_{1}}\colon j=1,\ldots,N\}$, where $\bar{Y}^{j}$ and $V^{j}$ are the maximum likelihood estimators of $\mu_{j}$ and $\sigma_{j}^{2}$, and keep the populations with the largest statistics for inference. We will not analyze any specific choice of selection criterion because, as we shall see, all of them are operationally identical under the considered form of inference. Thus, without loss of generality, suppose that $\mu_{j}$ is selected for inference if $S_{j}>T_{n_{1},N}(S_{-j})\equiv T_{j}$, where $S_{-j}$ contains all $t$-statistics except $S_{j}$. This threshold could be $T_{j}=\max(S_{-j})$, or it could be a constant significance threshold obtained from a formal hypothesis test, possibly with a multiplicity correction if $N$ is large. Furthermore, assume that, from each of the selected populations, we collect $n_{2}$ extra samples to increase inferential power: $Y_{n_{1}+1}^{j},\ldots,Y_{n}^{j}$, where $n=n_{1}+n_{2}$ is the total sample size. This is not necessary for the discussion, but mimics standard procedure in many applications.
@@
In problems of this type, it is common to conduct inference on each of the selected means conditionally on the data from the other $N-1$ populations, in addition to conditioning on the selection event, so as to eliminate the $2(N-1)$ nuisance parameters. For example, let us assume that $\mu_{1}$ is one of the selected means. Selective inference on $\mu_{1}$ is then to be conducted from the model $Y_{1}^{1},\ldots,Y_{n}^{1}\sim N(\mu_{1},\sigma_{1}^{2})$ with selection event $S_{1}>t_{1}=T_{1}(s_{2},\ldots,s_{N})$, which only depends on the remaining populations through the (now constant) threshold $t_{1}$. In a selective many-parameter problem, such reduction is often possible and enables analysis of the high-dimensional model in terms of a few independent problems involving low-dimensional selected parameters. Therefore, in our empirical investigations, we shall simply consider selection models of the form $Y_{1},\ldots,Y_{n}\sim N(\mu,\sigma^{2})$ with selection event $S>t$, for some fixed $t\in\mathbb{R}$, where the selected population index has been dropped. The parameter is $\theta=(\mu,\sigma^{2})$.
Let $(\bar{Y}_{n},V_{n})$ be the maximum likelihood estimator of $\theta$. We have

### v3

@@ 3 Asymptotic expansion of the likelihood
3.1 Main result
First, we examine how selection acts on Gaussian models. Let $Z\sim N(h,\Sigma)$, with parameter $h\in\mathbb{R}^{p}$ and fixed positive definite covariance $\Sigma\in\mathbb{R}^{p\times p}$, and consider the selection model induced by an arbitrary selection function $p(z)$. The selective log-likelihood ratio for this model around $h=0$ is given by
$\log\frac{f_{h}(Z)}{f_{0}(Z)}=h^{\top}\Sigma^{-1}Z-\frac{1}{2}h^{\top}\Sigma^{-1}h+\log\frac{\varphi(0)}{\varphi(h)},\quad\varphi(h)=E_{h}[p(Z)].$ (6)
@@ 3.1 Main result
The main result of this paper, Theorem 1, establishes conditions under which the sequence of selective log-likelihood ratios of a given selection model $\mathcal{M}_{n}(u_{n})$ admits an asymptotic expansion of the form (6) for a certain specification of selection function, which in general depends on the true generating parameter $\theta$. It is worth remarking that, since selection functions can vary arbitrarily with $n$, $\{\mathcal{M}_{n}(u_{n})\colon n\in\mathbb{N}\}$ does not in general converge to a fixed Gaussian selection model. Instead, it can, under suitable conditions, be approximated by a sequence of such.
Before stating the Theorem, we need to introduce some notation.
Notation. In the non-selective model $\mathcal{M}_{n}$, let $l_{\theta}(y_{i})=\log f_{\theta}(y_{i})$ denote the log-likelihood for the $i$-th observation, $\triangledown l_{\theta}(y_{i})$ the score, and $I_{\theta}=E_{\theta}[\triangledown l_{\theta}(Y_{i})\triangledown l_{\theta}(Y_{i})^{\top}]$ the per-observation Fisher information, assumed to be positive definite for all $\theta\in\Theta$. For a function $g:I\subseteq\mathbb{R}^{p}\to\mathbb{R}$, consider its Lipschitz norm,
@@
where $\|\cdot\|$ is the Euclidean norm; its supremum norm, $\|g\|_{\infty}=\sup\{|g(x)|\colon x\in I\}$; and its bounded-Lipschitz norm, $\|g\|_{\text{BL}}=\|g\|_{\infty}+\|g\|_{\text{L}}$. For a $J\subseteq I$, let $g_{|J}$ be the restriction of $g$ to $J$. The bounded-Lipschitz distance between the distributions of two $p$-dimensional random vectors $X$ and $Y$ is
@@
Suppose that $\{F_{\theta}\colon\theta\in\Theta\}$ is DQM, define $Z_{n}\equiv Z_{n}(\theta)=n^{-1/2}I_{\theta}^{-1}\sum_{i=1}^{n}\triangledown l_{\theta}(Y_{i})$, and assume that the sequence of selection functions is uniformly bounded:
@@
For a fixed $\theta\in\Theta$, assume that $M(x)=E_{\theta}[\exp\{x^{\top}\triangledown l_{\theta}(Y_{1})\}]<\infty$ in a neighborhood of the origin, that there exists a sequence of measurable functions $\{p^{*}_{n}(\cdot;\theta)\colon\mathbb{R}^{p}\to[0,\infty)\colon n\in\mathbb{N}\}$ satisfying $p^{*}_{n}(z;\theta)=E_{\theta}\left[p_{n}(Y^{n})\mid Z_{n}=z\right]$ for all $z$ in the support of $Z_{n}$, and that either of the following conditions are met:
@@
$\log\frac{f_{\theta+h_{n}/\sqrt{n}}(Y^{n}\mid u_{n})}{f_{\theta}(Y^{n}\mid u_{n})}=h^{\top}I_{\theta}Z_{n}-\frac{1}{2}h^{\top}I_{\theta}h+\log\frac{\varphi_{n}^{*}(0)}{\varphi_{n}^{*}(h)}+o_{F^{u_{n}}_{\theta}}(1).$ (11)
@@
An asymptotic connection is therefore established between the original model and a Gaussian selection model with base distribution $N(h,I_{\theta}^{-1})$ and selection function $p_{n}^{*}(z;\theta)$. In particular, the result implies that, to first order, the selective likelihood depends on the data only through the normalized score $Z_{n}$, so $Z_{n}$ is asymptotically sufficient and provides a natural asymptotic pivot for inference. Testing applications are explored in later sections.
The proof of Theorem 1 can be found in the supplementary material along with the proofs of all the other theoretical results of the paper. It takes as starting point the LAN property of the underlying, non-selective model, from which the first two terms of the expansion can be derived. The main technical difficulty lies in validating the approximation of the probability ratios $\varphi_{n}(\theta+n^{-1/2}h)/\varphi_{n}(\theta)$ by their Gaussian counterparts. This requires consideration of the moment generating function of $Z_{n}$ given selection. The existence of $M(x)$, together with the regularity assumptions on the selection functions in the limiting Gaussian models, prove to be sufficient to ensure uniform validity. While the regularity conditions on the selection functions are relatively mild, they do exclude scenarios where the probability of selection vanishes asymptotically. In such cases, the Gaussian approximation may fail to capture the behavior of the posterior or likelihood ratios, and the theoretical guarantees of Theorem 1 no longer apply. We therefore add a note of caution that the current framework is restricted to selection events with probabilities bounded away from zero.
[…]
3.2 Vanishing selection probabilities
The requirement that $\varphi_{n}(\theta)$ is bounded away from zero is needed to ensure that the Gaussian approximation of the normalized score remains valid after conditioning on selection and to preserve the scale of likelihood ratios. This condition guarantees asymptotic regularity of the model even if selection is deterministic, that is, when the selection functions are of the form $p_{n}(y^{n})=\mathbf{1}(y^{n}\in A_{n})$ for some events $A_{n}$. However, it is well known that, in these scenarios, if $y^{n}$ falls close to the boundary of $A_{n}$, irregularities can arise, leading to poor inferential performance unless $n$ is very large.
Scenarios with $\varphi_{n}(\theta)\to 0$ can arise in practice, particularly in models involving many parameters. A clear example arises when inferring the mean of a Gaussian model when the true mean is negative and the selection region is $A=(t_{n},\infty)$ for some sequence $t_{n}\geq 0$. Extending the framework to handle such vanishing-probability cases is an important direction for future work. While this asymptotic regime is not covered by the Theorem, in some scenarios it is possible to verify the boundedness condition either exactly or asymptotically:
@@ 3.2 Vanishing selection probabilities
    If the selection condition is rejection of a simple hypothesis $H_{0}\colon\theta=\theta_{0}$ at a level $\alpha$, then $\varphi_{n}(\theta)\geq\alpha$ for all $n$, provided the test is unbiased. Thus, as long as $\alpha$ does vanish as $n\to\infty$, the condition is satisfied. This is typically the case for multiple comparisons models, such as the one described in Section 3.5, with a bounded number of parameters.
@@
Selection rules that do not allow for the existence of $\sqrt{n}$-consistent estimators are typically deterministic and yield inferences with very limited power. Their lack of power often requires the collection of additional data to draw meaningful conclusions about the parameter, leading to a data carving scenario (Example 3.2) in which estimation of $\theta$ with the required accuracy is feasible.
3.3 Examples of common selection mechanisms
The following examples are representative of the main selection mechanisms encountered in selective inference.
@@ 3.3 Examples of common selection mechanisms
Deterministic selection. Let $Y_{1},\ldots,Y_{n}$ be an IID sample from an exponential family distribution with density
$f_{\theta}(y_{i})=h(y_{i})\exp\left\{\theta^{\top}y_{i}-A(\theta)\right\},\quad\theta\in\Theta.$ (12)
@@
Data carving. Consider the same setting as in the previous example, but with selection condition given by $\bar{Y}^{(1)}_{n}\in E_{n}$, where $\bar{Y}^{(1)}_{n}$ is the mean of the first $n/2$ samples, where $n$ is assumed to be even. Then,
@@
Condition 2 of the Theorem is satisfied provided $\sup\{\|\log f_{\bar{Y}^{(1)}_{n}}\|_{\text{L}}\colon n\geq 1\}<\infty$. Conditioning on selection based on a subsample of the data was introduced by Fithian et al. (2017) and is commonly referred to as data carving. The main operational consequence of data carving is a smoothing of the selection function.
@@
Randomization. Consider the same setting, but suppose now that the selection mechanism acts on $U_{n}=\sqrt{n}\bar{Y}_{n}+W$, where the noise $W$ has a known density $f_{W}(w)$, so that selection takes the form $U_{n}\in E_{n}$. Then, $p^{*}_{n}(z;\theta)=P\{\sqrt{n}\bar{y}_{n}(z)+W\in E_{n}\}$, so
@@
which is bounded provided $f_{W}(w)$ is bounded. Proposed by Tian and Taylor (2018), randomized selection has become common in practice in post-selection inference, as it enables more powerful analyses. Similarly to data carving, the randomization here tends to smooth the selection function
@@
Condition on randomized statistic. Consider the same setting as in the previous example. In cases where knowledge of the precise form of $E_{n}$ is unavailable to the statistician, or too complex to be conditioned on, it is common to condition on the observed value $U_{n}$ instead (García Rasines and Young, 2023; Leiner et al., 2023; Neufeld et al., 2024). In that case, the selection function becomes $p^{*}_{n}(z;\theta)=f_{W}\{u_{n}-\triangledown^{2}A(\theta)z+\sqrt{n}\triangledown A(\theta)\}$, which satisfies the required regularity conditions for most standard noise distributions, such as the standard Gaussian or Laplace distributions employed by Tian and Taylor (2018).
3.4 Exponential model (ctd.)
We will illustrate the four previous types of selection mechanism in the simple model of Example 2.1. A deterministic selection mechanism would be of the form $\bar{Y}_{n}<t_{n}$ for some predetermined threshold value $t_{n}$, which might have been obtained through a power analysis or in some other data-independent way. For a given $\theta$ and realization $\bar{Y}_{n}=\bar{y}_{n}$, the normalized score for this model is $z=\theta^{2}\sqrt{n}(1/\theta-\bar{y}_{n})$, and the selection function in the asymptotic model becomes $p^{*}_{n}(z;\theta)=\mathbf{1}\{z>\theta^{2}\sqrt{n}(1/\theta-t_{n})\}$. If the analogous selection rule is applied only to half the samples, so that inference is performed whenever $\bar{Y}_{n}^{(1)}<t_{n}$ (data carving), then, assuming $n$ is even and noticing that $Z_{n}$ follows a shifted and scaled gamma distribution, we obtain, after some manipulation,
@@
Finally, suppose selection acts on $U_{n}=\sqrt{n}\bar{Y}_{n}+W$, where $W\sim N(0,\sigma^{2}_{W})$ for some pre-determined $\sigma_{W}>0$, so that the sample gets selected if $U_{n}<t_{n}$ (randomization). Then,
$p^{*}_{n}(z;\theta)=\Phi\left\{\frac{1}{\sigma_{W}}\left(t_{n}-\frac{\sqrt{n}}{\theta}+\frac{z}{\theta^{2}}\right)\right\}.$
If, instead, we condition on the observed value of the randomized statistic $U_{n}$, $u_{n}$ (perhaps due to unavailability of knowledge of the precise threshold $t_{n}$ applied), the selection function becomes
$p^{*}_{n}(z;\theta)=\frac{1}{\sigma_{W}}\phi\left\{\frac{1}{\sigma_{W}}\left(u_{n}-\frac{\sqrt{n}}{\theta}+\frac{z}{\theta^{2}}\right)\right\}.$
Figure 2 shows the functions $r_{n}(h;\theta)=\log\{\varphi_{n}(\theta+h/\sqrt{n})/\varphi_{n}(\theta)\}$ and $r^{*}_{n}(h;\theta)=\log\{\varphi^{*}_{n}(h)/\varphi^{*}_{n}(0)\}$ in the four situations considered above, for $\theta=2$, $n=40$, and $h\in[-2,2]$. The main step in the proof of Theorem 1 is to establish that the two functions approximate each other uniformly in compact sets of $h$ as $n\to\infty$. This figure provides visual evidence of the approximation across the range of values of $h$ considered. In all cases, the functions $\varphi^{*}_{n}(h)$ where computed by numerical integration of $I_{\theta}^{1/2}\phi(I_{\theta}^{1/2}z)p^{*}_{n}(z;\theta)$ over $z$. We have seen in Figure 1 how accuracy of the approximation translates to closeness of inferences derived in the true exponential and asymptotically equivalent Gaussian selection models.
@@ 3.5 Inference on winners
In this section we explore a classic example of selection bias, where inference is conducted on the most promising parameters chosen from a large set of candidates. This constitutes a paradigmatic application of selective inference in the IID setting. In such problems, $N$ independent populations under investigation, with respective parameters $\theta_{1},\ldots,\theta_{N}$, are ranked based on a data-dependent criterion (typically some measure of significance of the parameter), and only the top-ranked populations are selected for further analysis. For example, we may compute NN $p$-values of significance and select for formal inference only those parameters whose respective $p$-values are among the smallest $K<N$ or fall below a fixed significance threshold $t$. Once a subset of parameters has been selected, further data may be collected on the selected populations, leading to a data-carving scenario, as described in the previous section. This type of situation is common in clinical studies, where multiple treatments are tested and only the most promising ones are advanced for further investigation. Another common application occurs in genomics, where a vast number of gene expressions are measured, and only those showing the strongest signals, such as significant associations with a disease, are formally analyzed.
@@
We will consider the following specification of the problem. For $j\in\{1,\ldots,N\}$ and $i\in\{1,\ldots,n_{1}\}$, let $Y_{i}^{j}\sim N(\mu_{j},\sigma^{2}_{j})$ independently, where all $\mu_{j}$’s and $\sigma^{2}_{j}$’s are unknown. That is, we have random samples of size $n_{1}$ from $N$ different Gaussian populations. Suppose that we are only interested in means with large values. Accordingly, we compute NN $t$-statistics, $\{S_{j}=\bar{Y}^{j}/\sqrt{V^{j}/n_{1}}\colon j=1,\ldots,N\}$, where $\bar{Y}^{j}$ and $V^{j}$ are the maximum likelihood estimators of $\mu_{j}$ and $\sigma_{j}^{2}$, and keep the populations with the largest statistics for inference. We will not analyze any specific choice of selection criterion because, as we shall see, all of them are operationally identical under the considered form of inference. Thus, without loss of generality, suppose that $\mu_{j}$ is selected for inference if $S_{j}>T_{n_{1},N}(S_{-j})\equiv T_{j}$, where $S_{-j}$ contains all $t$-statistics except $S_{j}$. This threshold could be $T_{j}=\max(S_{-j})$, or it could be a constant significance threshold obtained from a formal hypothesis test, possibly with a multiplicity correction if $N$ is large. Furthermore, assume that, from each of the selected populations, we collect $n_{2}$ extra samples to increase inferential power: $Y_{n_{1}+1}^{j},\ldots,Y_{n}^{j}$, where $n=n_{1}+n_{2}$ is the total sample size. This is not necessary for the discussion, but mimics standard procedure in many applications.
@@
In this type of problems, it is common to conduct inference on each of the selected means conditionally on the data from the other N1N-1 populations, in addition to conditioning on the selection event, so as to eliminate the 2(N1)2(N-1) nuisance parameters. For example, let us assume that $\mu_{1}$ is one of the selected means. Selective inference on $\mu_{1}$ is then to be conducted from the model $Y_{1}^{1},\ldots,Y_{n}^{1}\sim N(\mu_{1},\sigma_{1}^{2})$ with selection event $S_{1}>t_{1}=T_{1}(s_{2},\ldots,s_{N})$, which only depends on the remaining populations through the (now constant) threshold $t_{1}$. In a selective many-parameter problem, such reduction is often possible and enables analysis of the high-dimensional model in terms of a few independent problems involving low-dimensional selected parameters. Therefore, in our empirical investigations, we shall simply consider selection models of the form $Y_{1},\ldots,Y_{n}\sim N(\mu,\sigma^{2})$ with selection event $S>t$, for some fixed $t\in\mathbb{R}$, where the selected population index has been dropped.
Let $(\bar{Y}_{n},V_{n})$ be the maximum likelihood estimator of $\theta=(\mu,\sigma^{2})$. We have